OpenStudy (anonymous):

Solve (sqrt x + 5) - (sqrt x + 7) = -10

OpenStudy (anonymous):

can u help?? @nickersia ?

OpenStudy (anonymous):

if i help u than u have to help me

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

√(x) + 5 = -7 Isolate x by first subtracting 5 from each sides √(x) + 5 = -7 - 5 = - 5 √(x) (+ 0)= -12 √(x) = -12 To get rid of a square root sign you would have to square it. And you need to do this to both sides (say if x was 4 the square root would be 2 and squared again would be 4) √(x)^2 = (-12)^2 Since the square root and square cancel out you can simplify x = (-12)^2 Negative times a negative is postive and 12 times 12 is 144 x = 144 Hope this helps :) ****BTW. I noticed from the other person's answer that this can be interpreted two ways since you neglected to include parentheses. I assumed it was √(x) + 5 = -7 If it is √(x + 5) = -7 You would again square both sides √(x + 5)^2 = (-7) ^2 Simplify x+5=49 Subtract 5 x=44

OpenStudy (anonymous):

44

OpenStudy (anonymous):

so now help me

OpenStudy (anonymous):

ok sure what is the question?

OpenStudy (anonymous):

wait i saw the you only took the first part of my question the whole question is this \[\sqrt{x + 5} - \sqrt{x + 7} = -10\]

OpenStudy (anonymous):

Helloo??/

OpenStudy (anonymous):

@campbell_st can u help??

OpenStudy (campbell_st):

sorry the question doesn't seem to make a lot of sense to me, just wondering if there is something missing....

OpenStudy (anonymous):

the question is this solve \[\sqrt{x + 5} - \sqrt{x + 7} = -10\]

OpenStudy (anonymous):

according to wolframalpha it has no solutions... xD and, for example if you make it like: sqrt(x+5)+10=sqrt(x+7) and then power it, you'll have x+5+20sqrt(x+5)+100=x+7 sqrt(x+5)=-98/20 , so you'll be in the complex and i don't know if there's a solution over there

OpenStudy (anonymous):

ok